# Geometry of the moduli space of Hermitian-Einstein connections on manifolds with a dilaton

### v1

@@ Abstract
We demonstrate that the moduli space of Hermitian-Einstein connections $\mathscr{M}^{*}_{\mathrm{HE}}(M^{2n})$ of vector bundles over compact non-Gauduchon Hermitian manifolds $(M^{2n},g,\omega)$ that exhibit a dilaton field $\Phi$ admit a strong Kähler with torsion structure provided a certain condition is imposed on their Lee form $\theta$ and the dilaton. We find that the geometries that satisfy this condition include those that solve the string field equations or equivalently the gradient flow soliton type of equations. In addition, we demonstrate that if the underlying manifold $(M^{2n},g,\omega)$ admits a holomorphic and Killing vector field $X$ that leaves $\Phi$ also invariant, then the moduli spaces $\mathscr{M}^{*}_{\mathrm{HE}}(M^{2n})$ admits an induced holomorphic and Killing vector field $\alpha_{X}$. Furthermore, if $X$ is covariantly constant with respect to the compatible connection $\widehat{\nabla}$ with torsion a 3-form on $(M^{2n},g,\omega)$, then $\alpha_{X}$ is also covariantly constant with respect to the compatible connection $\widehat{\mathcal{D}}$ with torsion a 3-form on $\mathscr{M}^{*}_{\mathrm{HE}}(M^{2n})$ provided that $K^{\flat}\wedge X^{\flat}$ is a $(1,1)$-form with $K^{\flat}=\theta+2d\Phi$ and $\Phi$ is invariant under $X$ and $IX$, where $I$ is the complex structure of $M^{2n}$.
@@ 1 Introduction
Hermitian-Einstein connections on four-dimensional manifolds are special because such connections satisfy the instanton anti-self-duality condition. The anti-self-duality condition depends on the choice of orientation on the underlying 4-dimensional manifold $M^{4}$ and not on the choice of complex structure on $M^{4}$. This is provided that the choice of complex structure is compatible with the orientation chosen to impose the anti-self-duality condition. As a result the moduli spaces of instantons $\mathscr{M}_{\mathrm{asd}}^{*}(M^{4})$ can exhibit “exotic” geometric structures after making judicious choices of a complex structure on $M^{4}$. In particular, if $M^{4}$ admits a hyper-Kähler, (oriented) bi-KT, hyper-Kähler with torsion (HKT) or (oriented) bi-HKT structure [^3] Bi-KT and bi-HKT structures are also known as generalised Kähler and generalised hyper-Kähler structures [24, 25], respectively., then $\mathscr{M}_{\mathrm{asd}}^{*}(M^{4})$ exhibits the same structure as well [6, 24, 26]– in all cases the torsion 3-form [^4] All Hermitian manifolds $(M^{2n},g,\omega)$ admit a unique compatible connection $\widehat{\nabla}$ with skew-symmetric torsion $H$, where $H=-\iota_{I}d\omega$, ^$\widehat{\nabla}_{i}Y^{j}=D_{i}Y^{j}+\frac{1}{2}H^{j}{}_{ik}Y^{k}$ and $D$ is the Levi-Civita connection. $\mathcal{H}$ of the moduli space either vanishes or it is a closed 3-form, $d\mathcal{H}=0$ i.e. the KT, bi-KT, HKT and bi-HKT structures are strong.
@@
The choice of the Gauduchon gauge to investigate the geometry of $\mathscr{M}_{\mathrm{HE}}^{*}(M^{2n})$ is not much of a restriction on the geometry of the underlying compact manifold $M^{2n}$ as every Hermitian metric in its conformal class admits a Gauduchon metric. Moreover, the Hermitian-Einstein condition on a connection $A$ of a vector bundle $E$, especially with gauge group $SU(r)$, depends on the conformal class $[g]$ of the metric $g$ instead of $g$ itself. The same applies for the conditions that the tangent vectors of the moduli space $\mathscr{M}^{*}_{\mathrm{HE}}$ obey, see section 2.1. So provided that the geometry of $M^{2n}$ is not subject to some other constraint, one can always choose the Gauduchon metric to perform the investigation of the moduli space geometry. However, if another condition is imposed on the geometry of $M^{2n}$, this will not be the case whenever the additional condition is not preserved upon the rescalling $g\rightarrow e^{\Lambda}g$ of the metric. Such a rescalling is required to bring a Hermitian metric into the Gauduchon gauge. For example, if $(M^{2n},g,\omega)$ admits a strong KT structure, as $H=-\iota_{I}d\omega$, the closure $dH=0$ of the 3-form torsion $H$ is not preserved upon an arbitrary rescalling of the metric and so of $\omega\rightarrow e^{\Lambda}\omega$. The same applies whenever the holonomy of the connection with torsion $\widehat{\nabla}$ is restricted to lie in a proper subgroup of $U(n)$ or $(g,H)$ satisfy an additional condition that depends on $g$ and not its conformal class $[g]$. These kind of additional restrictions on the Hermitian geometry of $M^{2n}$ are common place in the context of various applications in physics and in geometry. Furthermore, even though the conditions that the tangent vector of $\mathscr{M}_{\mathrm{HE}}^{*}(M^{2n})$ satisfy depend only on the conformal class $[g]$ of $g$, the geometry of $\mathscr{M}_{\mathrm{HE}}^{*}(M^{2n})$, like the metric and Hermitian form, depends on $g$ itself instead of its conformal class. As a result different choices of $g$ lead to different geometries on $\mathscr{M}_{\mathrm{HE}}^{*}(M^{2n})$.
@@
$(a_{1},a_{2})_{\mathscr{A}^{*}}=\int_{M}d^{2n}x\,\sqrt{g}\,\,e^{-2\Phi}\,\,\langle a_{1},a_{2}\rangle_{\mathfrak{g}}\leavevmode\nobreak\ ,$ (1.1)
@@
$D^{i}\big{(}e^{-2\Phi}K^{\flat}_{i}\big{)}=0\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ K^{\flat}\equiv\theta+2d\Phi\leavevmode\nobreak\ ,$ (1.2)
where $K^{\flat}$ is the 1-form associated to the vector field $K$, i.e. $K^{\flat}(Y)=g(K,Y)$. This condition is the Gauduchon condition for another metric in the conformal class $[g]$ of $g$. But for a prescribed metric $g$ on $M^{2n}$, i.e. a metric that satisfies a condition on $M^{2n}$ that depends on $g$ and not its conformal class, (1.2) is a restriction on the geometry of $M^{2n}$. This condition is satisfied by a wide range of manifolds. These include those that occur in heterotic string supersymmetric compactifications [7, 8] and those that serve as target spaces for conformally invariant sigma models with metric and $B$-field couplings, i.e. those that satisfy the string field equations [30, 31, 32, 33, 34]. In the related context of Hermitian flows, for a review see [35] and references therein, we shall demonstrate that compact gradient flow solitons solve (1.2).
@@
The second theorem concerns the properties of the vector fields $\alpha_{X}$ induced by holomorphic and $\widehat{\nabla}$-covariantly constant vector fields $X$ on $M^{2n}$. Under the same assumptions on $(M^{2n},g,\omega)$ as in the previous theorem, we have the following.
@@
Let $X$ be $\widehat{\nabla}$-parallel and holomorphic vector field on $(M^{2n},g,\omega)$, and $X\Phi=Y\Phi=0$ with $Y=-IX$, where $I$ is the complex structure of $M^{2n}$. Then, $\alpha_{X}$ is $\widehat{\mathcal{D}}$-covariantly constant on $\mathscr{M}^{*}_{\mathrm{HE}}(M^{2n})$ provided that $X^{\flat}\wedge K^{\flat}$ is a $(1,1)$-form on $M^{2n}$, where $K^{\flat}=\theta+2d\Phi$ and $\widehat{\mathcal{D}}$ is the compatible connection on $\mathscr{M}^{*}_{\mathrm{HE}}(M^{2n})$ with skew-symmetric torsion.
@@
$\widehat{R}_{ij}-\frac{1}{4}P_{ij}=2\widehat{\nabla}_{i}X^{\flat}_{j}\leavevmode\nobreak\ ,$ (1.3)
@@
There are several options to choosing $P$ and $dH$. First, for $P=0$ and $H$ closed, $dH=0$, this is the 1-loop scale invariance condition for bosonic sigma models with couplings the metric and the $B$-field, $H=dB$ or equivalently it is the condition for steady flow solitons [36, 37, 38, 39].
@@
$P_{ij}=-\frac{\alpha^{\prime}}{4}\big{(}\tilde{R}_{ikmn}\tilde{R}_{j}{}^{kmn}-F(C)_{ik\mathrm{a}\mathrm{b}}F_{j}{}^{k\mathrm{a}\mathrm{b}}\big{)}\leavevmode\nobreak\ ,$ (1.4)
@@
$dH=\frac{\alpha^{\prime}}{4}\big{(}\tilde{R}^{i}{}_{j}\wedge\tilde{R}^{j}{}_{i}-F(C)^{\mathrm{a}}{}_{\mathrm{b}}\wedge F(C)^{\mathrm{b}}{}_{\mathrm{a}}\big{)}\leavevmode\nobreak\ ,$ (1.5)
@@
$\widehat{R}_{ij}-\frac{1}{4}P_{ij}+2\widehat{\nabla}_{i}\partial_{j}\Phi=0\leavevmode\nobreak\ ,$ (1.6)
@@
The relationship between (1.3) and (1.6) has been explored from different perspectives. From the perspective of scale vs conformal invariance, Polchinski [43] demonstrated that under certain conditions, which include unitarity and a discrete spectrum of operator scaling dimensions, that two-dimensional scale invariant Quantum Field Theories are actually conformally invariant. In the context of two-dimensional sigma models $P=dH=0$, the relations between (1.3) and (1.6) has been investigated by Hull and Townsend in [44]. In the same context, it has been demonstrated in [23] via a Perelman style of argument [45] that if the sigma model target space is compact, then (1.3) implies (1.6). Equivalently, steady flow solitons are in fact gradient flow solitons [35]. A similar result can be established for heterotic sigma models with compact target space [46], i.e. for sigma models with $P$ and $dH$ given in (1.4) and (1.5), respectively.
@@
^$\widehat{\nabla}_{i}V^{j}=0\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ V^{\flat}_{i}\equiv X^{\flat}_{j}+2\partial_{j}\Phi\leavevmode\nobreak\ ,$ (1.7)
i.e. the vector field $V$ is $\widehat{\nabla}$-covariantly constant. In turn this implies that $V$ is Killing vector field and $\iota_{V}H=dV^{\flat}$.
@@ 2.1.1 The space of connections
$(a,A)\qquad$ $\rightarrow A+a\leavevmode\nobreak\ ,$ (2.2)
@@
$a=d_{A}\epsilon\equiv d\epsilon+[A,\epsilon]_{\mathfrak{g}}\leavevmode\nobreak\ ,$ (2.3)
@@
$\mathcal{I}(a)=-\iota_{I}a\leavevmode\nobreak\ ,$ (2.4)
@@
$(a_{1},a_{2})_{\mathscr{A}}=\int_{M^{2n}}d^{d}x\leavevmode\nobreak\ \sqrt{g}g^{ij}\langle a_{1i},a_{2j}\rangle_{\mathfrak{g}}\leavevmode\nobreak\ .$ (2.5)
@@
$\Omega_{\mathscr{A}}(a_{1},a_{2})\equiv(a_{1},\mathcal{I}a_{2})_{\mathscr{A}}=\frac{1}{(n-1)!}\,\int_{M^{2n}}\omega^{n-1}\wedge\langle a_{1}\wedge a_{2}\rangle_{\mathfrak{g}}\leavevmode\nobreak\ .$ (2.6)
@@ 2.1.2 The tangent space of Hermitian-Einstein connections
$F^{2,0}=F^{0,2}=0\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \omega\llcorner F\equiv\frac{1}{2}\omega^{ij}F_{ij}=\lambda\leavevmode\nobreak\ ,$ (2.7)
@@
$d_{A}a^{2,0}=d_{A}a^{0,2}=0\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \omega\llcorner d_{A}a=0\leavevmode\nobreak\ ,$ (2.8)
@@ 2.1.3 The tangent space of the moduli space of Hermitian-Einstein connections
$\omega\llcorner d_{A}\iota_{I}a=0\Longrightarrow D_{A}^{i}a^{h}_{i}+\theta^{i}a^{h}_{i}=0\leavevmode\nobreak\ ,$ (2.10)
@@
$\mathcal{O}\epsilon=-\omega\llcorner d_{A}^{c}d_{A}\epsilon=(D_{A}^{i}D_{Ai}+\theta^{i}D_{Ai})\epsilon=-\omega\llcorner d_{A}^{c}a\leavevmode\nobreak\ .$ (2.12)
@@
$(\alpha_{1},\alpha_{2})_{\mathscr{M}^{*}_{\mathrm{HE}}}\equiv(a^{h}_{1},a^{h}_{2})_{\mathscr{A}^{*}_{\mathrm{HE}}}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \Omega_{\mathscr{M}^{*}_{\mathrm{HE}}}(\alpha_{1},\alpha_{2})\equiv\Omega_{\mathscr{A}^{*}_{\mathrm{HE}}}(a^{h}_{1},a^{h}_{2})\leavevmode\nobreak\ ,$ (2.14)

### v2

@@ Abstract
We demonstrate that the moduli space of Hermitian-Einstein connections $\mathscr{M}^{*}_{\mathrm{HE}}(M^{2n})$ of vector bundles over compact non-Gauduchon Hermitian manifolds $(M^{2n},g,\omega)$ that exhibit a dilaton field $\Phi$ admit a strong Kähler with torsion structure provided a certain condition is imposed on their Lee form $\theta$ and the dilaton. We find that the geometries that satisfy this condition include those that solve the string field equations or equivalently the (generalised) gradient flow soliton type of equations. In addition, we demonstrate that if the underlying manifold $(M^{2n},g,\omega)$ admits a holomorphic and Killing vector field $X$ that leaves $\Phi$ also invariant, then the moduli spaces $\mathscr{M}^{*}_{\mathrm{HE}}(M^{2n})$ admits an induced holomorphic and Killing vector field $\alpha_{X}$. Furthermore, if $X$ is covariantly constant with respect to the compatible connection $\widehat{\nabla}$ with torsion a 3-form on $(M^{2n},g,\omega)$, then $\alpha_{X}$ is also covariantly constant with respect to the compatible connection $\widehat{\mathcal{D}}$ with torsion a 3-form on $\mathscr{M}^{*}_{\mathrm{HE}}(M^{2n})$ provided that $K^{\flat}\wedge X^{\flat}$ is a $(1,1)$-form with $K^{\flat}=\theta+2d\Phi$ and $\Phi$ is invariant under both $X$ and $IX$, where $I$ is the complex structure of $M^{2n}$.
@@ 1 Introduction
Hermitian-Einstein connections on four-dimensional manifolds are special because such connections satisfy the instanton anti-self-duality condition. The anti-self-duality condition depends on the choice of orientation on the underlying 4-dimensional manifold $M^{4}$ and not on the choice of complex structure on $M^{4}$. This is provided that the choice of complex structure is compatible with the orientation used to impose the anti-self-duality condition. As a result the moduli spaces of instantons $\mathscr{M}_{\mathrm{asd}}^{*}(M^{4})$ can exhibit “exotic” geometric structures after making judicious choices of a complex structure on $M^{4}$. In particular, if $M^{4}$ admits a hyper-Kähler, (oriented) bi-KT, hyper-Kähler with torsion (HKT) or (oriented) bi-HKT structure [^3] Bi-KT and bi-HKT structures are also known as generalised Kähler and generalised hyper-Kähler structures [24, 25], respectively., then $\mathscr{M}_{\mathrm{asd}}^{*}(M^{4})$ exhibits the same structure as well [6, 24, 26]– in all cases the torsion 3-form [^4] All Hermitian manifolds $(M^{2n},g,\omega)$ admit a unique compatible connection $\widehat{\nabla}$ with skew-symmetric torsion $H$, where $H=-\iota_{I}d\omega$, $\widehat{\nabla}_{i}Y^{j}=D_{i}Y^{j}+\frac{1}{2}H^{j}{}_{ik}Y^{k}$ and $D$ is the Levi-Civita connection. $\mathcal{H}$ of the moduli space either vanishes or it is a closed 3-form, $d\mathcal{H}=0$, i.e. the KT, bi-KT, HKT and bi-HKT structures are strong.
@@
The choice of the Gauduchon gauge to investigate the geometry of $\mathscr{M}_{\mathrm{HE}}^{*}(M^{2n})$ is not much of a restriction on the geometry of the underlying compact manifold $M^{2n}$ as every Hermitian metric in its conformal class admits a Gauduchon metric. Moreover, the Hermitian-Einstein condition on a connection $A$ of a vector bundle $E$, especially with gauge group $SU(r)$, depends on the conformal class $[g]$ of the metric $g$ instead of $g$ itself. The same applies for the conditions that the tangent vectors of the moduli space $\mathscr{M}^{*}_{\mathrm{HE}}$ obey, see section 2.1. So provided that the geometry of $M^{2n}$ is not subject to some other constraint, one can always choose the Gauduchon metric to perform the investigation of the moduli space geometry. However, if another condition is imposed on the geometry of $M^{2n}$, this will not be the case whenever the additional condition is not preserved upon the rescaling $g\rightarrow e^{\Lambda}g$ of the metric. Such a rescaling is required to bring a Hermitian metric into the Gauduchon gauge. For example, if $(M^{2n},g,\omega)$ admits a strong KT structure, as $H=-\iota_{I}d\omega$, the closure $dH=0$ of the 3-form torsion $H$ is not preserved upon an arbitrary rescaling of the metric and so of $\omega\rightarrow e^{\Lambda}\omega$. The same applies whenever the holonomy of the connection with torsion $\widehat{\nabla}$ is restricted to lie in a proper subgroup of $U(n)$ or $(g,H)$ satisfy an additional condition that depends on $g$ and not its conformal class $[g]$. These kind of additional restrictions on the Hermitian geometry of $M^{2n}$ are common place in the context of various applications in physics and in geometry. Furthermore, even though the conditions that the tangent vectors of $\mathscr{M}_{\mathrm{HE}}^{*}(M^{2n})$ satisfy depend only on the conformal class $[g]$ of $g$, the geometry of $\mathscr{M}_{\mathrm{HE}}^{*}(M^{2n})$, like the metric and Hermitian form, depends on $g$ itself instead of its conformal class. As a result different choices of $g$ lead to different geometries on $\mathscr{M}_{\mathrm{HE}}^{*}(M^{2n})$.
@@
$(a_{1},a_{2})_{\mathscr{A}^{*}}=\int_{M}d^{2n}x\,\sqrt{g}\,\,e^{-2\Phi}\,\,\langle a_{1},a_{2}\rangle_{\mathfrak{g}}\penalty 10000\ ,$ (1.1)
@@
$D^{i}\big(e^{-2\Phi}K^{\flat}_{i}\big)=0\penalty 10000\ ,\penalty 10000\ \penalty 10000\ \penalty 10000\ K^{\flat}\equiv\theta+2d\Phi\penalty 10000\ ,$ (1.2)
where $K^{\flat}$ is the 1-form associated to the vector field $K$, i.e. $K^{\flat}(Y)=g(K,Y)$. This condition is the Gauduchon condition for another metric in the conformal class $[g]$ of $g$. But for a prescribed metric $g$ on $M^{2n}$, i.e. a metric that satisfies a condition on $M^{2n}$ that depends on $g$ and not its conformal class, (1.2) is a restriction on the geometry of $M^{2n}$. This condition is satisfied by a wide range of manifolds. These include those that occur in heterotic string supersymmetric compactifications [7, 8] and those that serve as target spaces for conformally invariant sigma models with metric and $B$-field couplings, i.e. those that satisfy the string field equations [30, 31, 32, 33, 34]. In the related context of Hermitian flows, for a review see [35] and references therein, we shall demonstrate that compact (generalised) gradient flow solitons solve (1.2).
@@
The second theorem concerns the properties of the vector fields $\alpha_{X}$ induced by holomorphic and $^\widehat{\nabla}$-covariantly constant vector fields $X$ on $M^{2n}$. Under the same assumptions on $(M^{2n},g,\omega)$ as in the previous theorem, we have the following.
@@
Let $X$ be $^\widehat{\nabla}$-parallel and holomorphic vector field on $(M^{2n},g,\omega)$, and $X\Phi=Y\Phi=0$ with Y=IXY=-IX, where $I$ is the complex structure of $M^{2n}$. Then, $\alpha_{X}$ is $\widehat{\mathcal{D}}-$covariantly constant on $\mathscr{M}^{*}_{\mathrm{HE}}(M^{2n})$ provided that $X^{\flat}\wedge K^{\flat}$ is a $(1,1)$-form on $M^{2n}$, where $K^{\flat}=\theta+2d\Phi$ and $\widehat{\mathcal{D}}$ is the compatible connection on $\mathscr{M}^{*}_{\mathrm{HE}}(M^{2n})$ with skew-symmetric torsion.
@@
$\widehat{R}_{ij}-\frac{1}{4}P_{ij}=2\widehat{\nabla}_{i}X^{\flat}_{j}\penalty 10000\ ,$ (1.3)
@@
There are several options to choosing $P$ and $dH$. First, for $P=0$ and $H$ closed, $dH=0$, this is the 1-loop scale invariance condition for bosonic sigma models with couplings the metric and the $B$-field, $H=dB$, or equivalently it is the condition for steady flow (generalised) solitons [37, 38, 39, 40].
@@
[…]
@@
[…]
@@
$\widehat{R}_{ij}-\frac{1}{4}P_{ij}+2\widehat{\nabla}_{i}\partial_{j}\Phi=0\penalty 10000\ ,$ (1.6)
@@
The relationship between (1.3) and (1.6) has been explored from different perspectives. From the perspective of scale vs conformal invariance, Polchinski [44] demonstrated that under certain conditions, which include unitarity and a discrete spectrum of operator scaling dimensions, that two-dimensional scale invariant Quantum Field Theories are actually conformally invariant. In the context of two-dimensional sigma models $P=dH=0$, the relations between (1.3) and (1.6) has been investigated by Hull and Townsend in [45]. In the same context, it has been demonstrated in [23] via a Perelman style of argument [46] that if the sigma model target space is compact, then (1.3) implies (1.6). Equivalently, steady flow (generalised) solitons are in fact gradient flow (generalised) solitons [35]. A similar result can be established for heterotic sigma models with compact target space [47], i.e. for sigma models with $P$ and $dH$ given in (1.4) and (1.5), respectively.
@@
$\widehat{\nabla}_{i}V^{j}=0\penalty 10000\ ,\penalty 10000\ \penalty 10000\ \penalty 10000\ V^{\flat}_{i}\equiv X^{\flat}_{j}+2\partial_{j}\Phi\penalty 10000\ ,$ (1.7)
i.e. the vector field $V$ is $^\widehat{\nabla}$-covariantly constant. In turn this implies that $V$ is Killing vector field and $\iota_{V}H=dV^{\flat}$.
@@ 2.1.1 The space of connections
$(a,A)\qquad$ $\rightarrow A+a\penalty 10000\ ,$ (2.2)
@@
$a=d_{A}\epsilon\equiv d\epsilon+[A,\epsilon]_{\mathfrak{g}}\penalty 10000\ ,$ (2.3)
@@
$\mathcal{I}(a)=-\iota_{I}a\penalty 10000\ ,$ (2.4)
@@
$(a_{1},a_{2})_{\mathscr{A}}=\int_{M^{2n}}d^{d}x\penalty 10000\ \sqrt{g}g^{ij}\langle a_{1i},a_{2j}\rangle_{\mathfrak{g}}\penalty 10000\ .$ (2.5)
@@
$\Omega_{\mathscr{A}}(a_{1},a_{2})\equiv(a_{1},\mathcal{I}a_{2})_{\mathscr{A}}=\frac{1}{(n-1)!}\,\int_{M^{2n}}\omega^{n-1}\wedge\langle a_{1}\wedge a_{2}\rangle_{\mathfrak{g}}\penalty 10000\ .$ (2.6)
@@ 2.1.2 The tangent space of Hermitian-Einstein connections
$F^{2,0}=F^{0,2}=0\penalty 10000\ ,\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \omega\llcorner F\equiv\frac{1}{2}\omega^{ij}F_{ij}=\lambda\penalty 10000\ ,$ (2.7)
@@
$d_{A}a^{2,0}=d_{A}a^{0,2}=0\penalty 10000\ ,\penalty 10000\ \penalty 10000\ \penalty 10000\ \omega\llcorner d_{A}a=0\penalty 10000\ ,$ (2.8)
@@ 2.1.3 The tangent space of the moduli space of Hermitian-Einstein connections
$\omega\llcorner d_{A}\iota_{I}a=0\Longrightarrow D_{A}^{i}a^{h}_{i}+\theta^{i}a^{h}_{i}=0\penalty 10000\ ,$ (2.10)
@@
$\mathcal{O}\epsilon=-\omega\llcorner d_{A}^{c}d_{A}\epsilon=(D_{A}^{i}D_{Ai}+\theta^{i}D_{Ai})\epsilon=-\omega\llcorner d_{A}^{c}a\penalty 10000\ .$ (2.12)
@@
$(\alpha_{1},\alpha_{2})_{\mathscr{M}^{*}_{\mathrm{HE}}}\equiv(a^{h}_{1},a^{h}_{2})_{\mathscr{A}^{*}_{\mathrm{HE}}}\penalty 10000\ ,\penalty 10000\ \penalty 10000\ \penalty 10000\ \Omega_{\mathscr{M}^{*}_{\mathrm{HE}}}(\alpha_{1},\alpha_{2})\equiv\Omega_{\mathscr{A}^{*}_{\mathrm{HE}}}(a^{h}_{1},a^{h}_{2})\penalty 10000\ ,$ (2.14)